OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (-3, 4) and (2, 1)? Check all that apply.

OpenStudy (texaschic101):

first find the slope using slope formula : slope (m) = (y2 - y1) / (x2 - x1) (-3,4) x1 = -3 and y1 = 4 (2,1) x2 = 2 and y2 = 1 now we sub slope(m) = (1 - 4) / (2 - (-3) slope(m) = -3/(2 + 3) slope(m) = -3/5 now we will use point slope form on both sets of points : point slope formula is : y - y1 = m(x - x1) (-3,4) -- x1 = -3 and y1 = 4 slope(m) = -3/5 now we sub y - 4 = -3/5(x - (-3) y - 4 = -3/5(x + 3)====> here is one (2,1) x1 = 2 and y1 = 1 slope(m) = -3/5 y - y1 = m(x - x1) now we sub y - 1 = -3/5(x - 2) ====> here is the other one your 2 point slope equations are : y - 4 = -3/5(x + 3) y - 1 = -3/5(x - 2)